# Rapid Parameter Estimation for Merging Massive Black Hole Binaries Using ODE-Based Generative Models

### v1

@@ 3.1 Network architecture
A key challenge in space-based gravitational wave parameter estimation lies in managing the high-dimensional nature of its data. To address this, it is essential to develop a feature extraction network capable of transforming the high-dimensional input data into a more manageable, low-dimensional representation [29, 32, 60]. This transformation is critical for adjusting the dimensionality of the final feature vector to match that of the parameter vector $\theta$, enabling the network to effectively parameterize the conditional vector field $(t,x,\theta)\rightarrow v_{t,x}(\theta)$ [33]. Achieving accurate parameterization of this conditional vector field is fundamental to the construction of an effective network structure.
@@
In recent years, DeResNet networks have been commonly used as the backbone network for feature extraction in ground-based gravitational wave parameter estimations or space-based gravitational wave parameter estimation [57, 32, 33, 53], and Gated Linear Unit(GLU)-conditioning [61] can be optionally used for information aggregation of $(\theta,t)$.
However, if information about the actual parameter $\theta$ is not properly handled or retained during the process of aggregating $(\theta,t)$ information using GLU, it could potentially lead to a loss of information about $\theta$. To encode the time $t$ more effectively and ensure that our feature extraction network captures its dynamic nature, we introduce a frequency vector $f$. This frequency vector $f$ adapts dynamically to changes in the parameters of the wave source, allowing for a more precise encoding of time $t$. To capture the periodic features of time, we employ sine and cosine functions for encoding. Our network architecture, referred to as TResMlp, is shown in Figure 4.
@@
Compared to the DeResNet model that use GLU-conditioning for aggregating information about $(\theta,t)$ [57, 32, 33, 53], our network architecture, TResMlp, demonstrates faster training speed and higher precision with the same hidden layers. More experiments and comparisons of related models are described in detail in Section 4.
@@ 3.2 Coordinate Transformation and Model Generalization Strategy
To simplify the dataset and reduce the difficulty of training, the coalescence times of MBHBs are set within $1\pm 0.02$ day in the training set (see Sec. 2). While the response function of detector to GW signals varies in time with a one-year period, thus in principle the model trained on this dataset can only be applied to MBHBs that merge within $1\pm 0.02$ day. While, thanks to the generalization strategy in this section, the application scope of the model can be extended to the whole mission lifetime.
In this paper, the time of coalescence (dubbed $t_{c}$) is defined relative to a reference time $t_{\rm ref}$. The complexity of space-based GW data partly lies in the annual modulation of detector response due to the orbital motion of detector constellation around the sun. In other words, the mapping between GW data and MBHB parameters varies with $t_{\rm ref}$. This significantly increases the complexity of the training set, which would severely damage the training efficiency and the performance of the model. To tackle this challenge, we exploit the symmetry (invariance) of the system, fix the training data around a certain $t_{\rm ref0}$, and then the trained model is generalized via our transformation rule to make inferences at any time. The idea behind this is a paradigm that combines physical information with neural networks, and methods belonging to this family, such as the group-equivariant neural posterior estimation [62, 60], have shown broad application prospects.
@@
We describe the transformation rule as follows. We define two coordinate systems, namely the Solar System Barycenter (SSB), in which physical quantities are indexed with “S”, and the Taiji detector reference frame, in which physical quantities are indexed with “T”. The parameters which vary with the coordinate systems are $\{t_{c},\lambda,\beta,\psi\}$, where $t_{c}$ should be interpreted as the time of arrival at the origin of the coordinate system (relative to $t_{\rm ref}$). Since the data in consideration last only for one day, for preliminary parameter estimation tasks, the location of MBHB remains stationary relative to the Taiji reference frame, and the mapping between source parameters $\bm{\theta}_{T}$ and TDI response is independent of $t_{\rm ref}$. This is precisely the symmetry based on which we can achieve model generalization. The $t_{\rm ref}$-dependent coordinate transformation from Taiji frame to SSB frame is denoted as $f_{t_{\rm ref}}:\bm{\theta}_{T}\rightarrow\bm{\theta}_{S,{t_{\rm ref}}}$ (see Ref. [63] for the expression of $f$). The parameters are firstly inferred with the model trained around $t_{\rm ref0}$, hence dubbed $\bm{\theta}_{S,t_{\rm ref0}}$, and the targeted parameters $\bm{\theta}_{S,t_{\rm ref}}$ can be obtained following two steps:
$\bm{\theta}_{T}=f_{t_{\rm ref0}}^{-1}\left(\bm{\theta}_{S,t_{\rm ref0}}\right),\quad\bm{\theta}_{S,t_{\rm ref}}=f_{t_{\rm ref}}\left(\bm{\theta}_{T}\right),$ (12)
or in terms of posterior distributions:
$p\left(\bm{\theta}_{S,t_{\rm ref}}|d\right)=p\left(\bm{\theta}_{S,t_{\rm ref0}}|d\right)J\left(f^{\ }_{t_{\rm ref0}}\right)J\left(f^{-1}_{t_{\rm ref}}\right),$ (13)
with $J\left(f^{\ }_{t_{\rm ref0}}\right)$ and $J\left(f^{-1}_{t_{\rm ref}}\right)$ being the Jacobian determinants corresponding to $f_{t_{\rm ref0}}$ and $f^{-1}_{t_{\rm ref}}$, respectively. Formula (13) looks similar to the transformations in normalizing flows, except that the mapping $f_{t_{\rm ref}}$ is constructed based on our understanding of the system, instead of training a neural network. Training a network to represent this transformation is unnecessary since we possess comprehensive knowledge of its form and can effortlessly compute it. As a result, The embedding of this physical information greatly simplifies training tasks.
Below, we describe how models can be combined with the model generalization strategy. Firstly, during the dataset generation, the merger times of MBHBs are limited to 0.02 days centered at a specific reference time (the 1st day). This significantly simplifies the dataset’s complexity and reduces the challenge of MBHB parameter estimation. Subsequently, we employ VPFMPE as the loss function for training CNFs. We utilize the novel network architecture TResMlp as the gravitational wave feature extractor, labeling this approach as VPFMPE-TResMlp.
In the inference phase, we use the model generalization strategy to map the reference time data for inference to the reference time data fixed during training. This allows the model to be trained on only one day of data but tested and applied to data spanning an entire year. We use the P-P plot in Figure 6 to demonstrate the importance of using a model generalization strategy.
@@ 4 Experiments
To ensure unbiased parameter estimation, we examine our algorithm on injected GW signals and noise in this section. For the parameter estimation of the signals, we employed the four models mentioned above: VPFMPE-TResMlp, FMPE-TResMlp, FMPE-DeResNet [33], and NPE-DeResNet [32]. We applied the model generalization strategy 3.2 to all four models by default. Otherwise, due to the complexity of the dataset, the models would lack inference capability.
Table 3: In the 850 simulations, we compared the correlation coefficients between the true values and the recovered values for the four models. In addition, in the last column of the table, we have included the training time for the three models.
$\mathcal{M}_{c}$ $q$ $\chi_{z1}$ $\chi_{z2}$ $d_{L}$ Training time
VPFMPE-TResMlp 0.9978 0.9790 0.9723 0.8163 0.9408 $\approx$ 2.0 days
FMPE-TResMlp 0.9978 0.9787 0.9719 0.8139 0.9400 $\approx$ 2.0 days
FMPE-DeResNet 0.9970 0.9762 0.9563 0.7174 0.9397 $\approx$ 2.3 days
NPE-DeResNet 0.9261 0.8647 0.6620 0.1681 0.8634 $\approx$ 6.5 days
To demonstrate the effectiveness of our algorithm and validate that the recovered probability distribution accurately represents the confidence we should have in the signal parameters, we conducted extensive testing on simulated signals and noise. We injected 850 waveforms drawn from the prior, along with residual instrumental and foreground confusion noise. We will take the median of the recovered values as the point estimate and calculate the correlation coefficient with the true values of the parameters. The correlation coefficient reflects the degree of linear correlation between the true value distribution and the estimated values. A coefficient value of 1 indicates perfect consistency between the estimated and true values. Note that the reference times of these testing signals are randomly distributed throughout the year, rather than being fixed to the 1st day.
Our results are presented in Table 3. In the final column of the Table 3, we have included the approximate training times for the three models. The VPFMPE-TResMlp model achieved the best results and had the fastest training time. For $t_{c}$, $\varphi_{c}$, $\iota$, $\beta$, $\lambda$, and $\psi$, due to the presence of multimodality (i.e. there is more than one peak in the posterior distributions), using the median as a point estimate is not applicable.
@@
To date, existing research has not yet succeeded in providing unbiased estimates of MBHBs across the entire parameter space. For example, Ref. [36] only focused on four parameters, and Ref. [32] have limited to positive spins. Also, we use machine learning techniques for complete parameter estimation of MBHBs and achieve comparable results to nested sampling. Therefore, in subsequent comparisons, we will only compare the results of VPFMPE-TResMlp to those of the MCMC method.
The aim of our ODE-based generative models for MBHBs is to reproduce the comparable posterior distribution as the MCMC method. Thus, by comparing the results of Bayesian inference on the same event, the performance of the model can be better assessed. In this case, we use Nested Sampling [64, 65, 66] as a Monte Carlo technique for comparison experiments, as shown in Figure 7.
An illustration demonstrating an example using the same prior probabilities outlined in Table 1 is provided in Figure 7, where a test injection is subjected to confusion noise on the 30th reference day. The injected GW signal originates from an MBHB with a total signal-to-noise ratio of 1543, and its parameters are detailed in Table 4. Compared to the standard nested sampling technique, our model exhibited results similar to the traditional nested sampling, demonstrating consistency in the sense of the degeneracy among extrinsic parameters 4. This consistency aligns with the crucial physical characteristics of MBHBs. Note that compared to nested sampling, our machine learning process always produces similar posterior distributions on $t_{c}$, $d_{L}$, $\iota$, $\beta$, $\lambda$, and $\psi$ 7.
We have achieved accurate estimates of the sky position parameters. This is most advantageous since accurate and reliable estimates of the sky position parameters (longitude, latitude) are useful in guiding the search for electromagnetic observatories.
We have collated the true values of the injections alongside the results generated by both the model and the nested sampling process in Table  4. This provides a concise overview of the performance of each method.
@@
A noteworthy phenomenon is that our machine learning pipeline produces a posterior similar to the nested sampling approach in terms of extrinsic parameters, and the posterior distributions of both converge in the same order of magnitude.
In the final part of the results, we explore the computational efficiency of the algorithm. Table 5 shows the speed of generating posterior samples using nested sampling compared to the VPFMPE method. It is important to note that we used nested sampling to search near the true intrinsic parameter values of the injected MBHB. In practice, the time consumed by nested sampling is significantly longer than the times listed in the table. Even under these conditions, our VPFMPE method remains two orders of magnitude faster than the nested sampling method.
@@ 5 Conclusions
In this study, we have demonstrated the inaugural application of ODE-based generative models in the field of space-based GW detection, taking the parameter estimation of merging MBHBs as a representative example. We recommend the use of ODE-based generative models, particularly the VPFMPE method, as the mainstream machine learning inference approach.
Our method provides comprehensive statistical inference for the 11-dimensional parameters of MBHBs, establishing a new benchmark in this area. This approach, as opposed to likelihood-based inference methods, can lead to both faster training and more accurate estimation.
Furthermore, we propose the integration of model generation strategy based on the symmetry of the system with ODE-based generative models. This enables us to simplify the training task, while still obtaining a model that is applicable to the entire mission period. Note that the coordinate transformation and ODE-based generative models are theoretically applicable to other space-based missions such as LISA.
It is also noteworthy that our method exhibits robustness against challenges such as confusion noise and the complexities of varying reference times, highlighting its reliability and potential for broader applications. The significant speed of our method, surpassing traditional techniques by orders of magnitude, positions it as a valuable tool for preprocessing in global fitting. Moreover, the precise and reliable estimation of sky location parameters (longitude, latitude) proves beneficial in guiding the search for electromagnetic observatories.
This acceleration in data processing and analysis not only streamlines workflows but also broadens the horizons for gravitational wave data exploration. Overall, these advancements herald a transformative era in space-based GW detection, fortifying the capabilities of missions like Taiji and bringing us closer to unraveling the mysteries of the universe.

### v3

@@ 3.1 Network architecture
A key challenge in space-based gravitational wave parameter estimation lies in managing the high dimensionality of data. To address this, it is essential to develop a feature extraction network capable of transforming the high-dimensional input data into a more manageable, low-dimensional representation [29, 32, 67]. This transformation is critical for adjusting the dimensionality of the final feature vector to match that of the parameter vector $\theta$, enabling the network to effectively parameterize the conditional vector field $(t,x,\theta)\rightarrow v_{t,x}(\theta)$ [34]. Achieving accurate parameterization of this conditional vector field is fundamental to the construction of an effective network structure.
@@
In recent years, DeResNet networks have been commonly used as the backbone network for feature extraction in ground-based or space-based gravitational wave parameter estimations [65, 32, 34, 61], and Gated Linear Unit(GLU)-conditioning [68] can be optionally used for information aggregation of $(\theta,t)$.
However, if information about the actual parameter $\theta$ is not properly handled or retained during the process of aggregating $(\theta,t)$ information using GLU, it could potentially lead to a loss of information about $\theta$. To encode the time $t$ more effectively, we introduce a frequency vector $f$. This frequency vector $f$ adapts dynamically to changes in the parameters of the wave source, allowing for a more precise encoding of time $t$. Our network architecture, referred to as TResMlp, is shown in Figure 3.
@@
Compared to the DeResNet model that uses GLU-conditioning for aggregating information about $(\theta,t)$ [65, 32, 34, 61], our network architecture, TResMlp, demonstrates faster training speed and higher precision with the same hidden layers. More experiments and comparisons of related models are described in detail in Section 4.
@@ 3.2 Coordinate Transformation and Model Generalization Strategy
To simplify the dataset and reduce the difficulty of training, in the training set, the coalescence times $t_{c}$ of MBHBs are set within a small interval $[1-0.01,1+0.01]$ day, see Sec. 2. In principle, the model trained on this dataset can only be applied to MBHBs that merge within this range. While, using the generalization strategy introduced in this section, the application scope of the trained model can be extended to the whole mission lifetime.
The complexity of space-based GW data partly lies in the annual variation of response function due to the orbital motion of detector constellation around the sun. In other words, the mapping between data and parameters varies with $t_{f}$, see Eq. (1). Since the time duration of data in consideration is much smaller than the time scale of orbital motion, $t_{f}$ can be well approximated by $t_{\rm ref}$, thus $t_{\rm ref}$ labels the position of detector in the orbit. The time-dependence of response function significantly increases the complexity of the training set, and severely damage the training efficiency and the performance of the model. To tackle this challenge, we exploit the symmetry (invariance) of the system, fix $t_{\rm ref}$ in the training data, and then the trained model is generalized via coordinate transformation to make inferences at any time.
@@
To explain the so called “symmetry”, we define two coordinate systems, namely the Solar System Barycenter (SSB) frame, in which parameters are indexed with subscript “S”, and a coordinate frame fixed to the Taiji detector, in which parameters are indexed with subscript “T”. Parameters which differ in these two coordinate systems are $\{\lambda,\beta,\psi\}$. As is explained, in the SSB frame, the mapping between parameters $\bm{\theta}_{S}$ and TDI response varies with $t_{\rm ref}$. While in the Taiji frame, due to the short duration of data, the motion of GW source relative to the detector can be neglected, thus the mapping between parameters $\bm{\theta}_{T}$ and TDI response is independent of $t_{\rm ref}$. Essentially, the former $t_{\rm ref}$-dependent mapping can be regarded as the combination of the latter $t_{\rm ref}$-independent mapping and a $t_{\rm ref}$-dependent coordinate transformation of parameters between the SSB frame and Taiji frame, and the analytical formula of this coordinate transformation is known, given the information of orbit. This is the fundamental principle behind training data simplification and model generalization.
Following Refs. [69, 70], we denote as $O_{1}(\lambda,\beta,\psi)$ the rotation matrix between the polarization basis in the GW source frame and the observer frame (Taiji or SSB), and $O_{2}(t_{\rm ref})$ the rotation matrix between the SSB frame and Taiji frame (see Eq. (6) and Eq. (9) of Ref. [69] for their expressions). Parameters $\{\lambda,\beta,\psi\}$ can be transformed between Taiji and SSB frames by solving [70]:
$[\cos\lambda_{T}\cos\beta_{T},\sin\lambda_{T}\cos\beta_{T},\sin\beta_{T}]$ $=O_{2}^{T}(t_{\rm ref})[\cos\lambda_{S}\cos\beta_{S},\sin\lambda_{S}\cos\beta_{S},\sin\beta_{S}],$
$O_{1}(\lambda_{T},\beta_{T},\psi_{T})$ $=O_{2}^{T}(t_{\rm ref})O_{1}(\lambda_{S},\beta_{S},\psi_{S}).$ (13)
For brevity, we further denote the $t_{\rm ref}$-dependent coordinate transformation from Taiji frame to SSB frame as $\mathcal{F}_{t_{\rm ref}}:\bm{\theta}_{T}\rightarrow\bm{\theta}_{S,{t_{\rm ref}}}$.
The strategy of model generalization is described as follows: During the training stage, the model is trained on a dataset with $t_{c}\in[t_{\rm ref0}-0.01,t_{\rm ref0}+0.01]$, where $t_{\rm ref0}\equiv 1$. When using the trained model to perform parameter estimation on the test data (generally $t_{\rm ref}\neq t_{\rm ref0}$), we firstly obtain the estimated parameters $\bm{\theta}_{S,t_{\rm ref0}}$. Obviously these parameters are not the desired ones. Then, the desired parameters $\bm{\theta}_{S,t_{\rm ref}}$ can be deduced following two steps:
$\bm{\theta}_{T}=\mathcal{F}_{t_{\rm ref0}}^{-1}\left(\bm{\theta}_{S,t_{\rm ref0}}\right),\quad\bm{\theta}_{S,t_{\rm ref}}=\mathcal{F}_{t_{\rm ref}}\left(\bm{\theta}_{T}\right).$ (14)
By performing these steps on each posterior sample, we finally obtain the target posterior distribution
$p\left(\bm{\theta}_{S,t_{\rm ref}}|d\right)=p\left(\bm{\theta}_{S,t_{\rm ref0}}|d\right)J\left(\mathcal{F}^{\ }_{t_{\rm ref0}}\right)J\left(\mathcal{F}^{-1}_{t_{\rm ref}}\right),$ (15)
with $J\left(\mathcal{F}^{\ }_{t_{\rm ref0}}\right)$ and $J\left(\mathcal{F}^{-1}_{t_{\rm ref}}\right)$ being the Jacobian determinants corresponding to $\mathcal{F}_{t_{\rm ref0}}$ and $\mathcal{F}^{-1}_{t_{\rm ref}}$, respectively, and $d$ stands for the data. Formula (15) looks similar to the transformations in normalizing flows, except that mapping $\mathcal{F}_{t_{\rm ref}}$ is constructed analytically, instead of training a neural network. As a result, our knowledge about the system has greatly simplified the training task. This idea represents a paradigm of combining physical information with neural networks. Methods belonging to this family, such as the group-equivariant neural posterior estimation [71, 67], have shown bright prospects. For future space-based missions, once the detectors’ motions are obtained through orbit determination, we can calculate the specific forms of matrices $O_{1}$ and $O_{2}$, thereby applying Eq. (3.2) - Eq. (15) to conduct the model generalization strategy.
Below, we describe how models can be combined with the model generalization strategy. We employ VPFMPE as the loss function for training CNFs. We utilize the network architecture TResMlp as the gravitational wave feature extractor, labeling this approach as VPFMPE-TResMlp. In the training phase we only use the 1st day’s data. During the inference phase, we employ the model generalization strategy to obtain the desired posterior distribution. This allows the model to be trained using the simplified dataset but tested and applied to data in the whole mission lifetime. We use the P-P plot in Figure 5 to demonstrate the importance of using a model generalization strategy.
3.3 Comparison of methods
In this section, we injected a substantial number of MBHB signals with noise to a comparison between the efficacy of FM with CNFs and NPE with DNFs (Recent advances in machine learning for parameter estimation in MBHBs). We use the three mentioned methods for FM with CNFs to compare them with NPE with DNFs. we employed the four models mentioned above: VPFMPE-TResMlp, FMPE-TResMlp, FMPE-DeResNet, and NPE-DeResNet [32]. We applied the model generalization strategy 3.2 to all four models by default. Otherwise, due to the complexity of the dataset, the models would lack inference capability.
To ensure the fairness of the training process, all models were trained under uniform conditions, involving 400 epochs, with each batch containing 14,336 samples. All models were trained using a single NVIDIA A800 GPU. In terms of optimization strategy, we used a combination of cosine annealing [72] and the Adam optimiser [73] to gradually reduce the learning rate to zero over the course of the training process. Our model training is based on a dataset consisting of 5 million MBHB signals. To prevent overfitting, we reserved 5% of the dataset for validation purposes. During the training process, no signs of overfitting were detected. During the testing phase of the model, we performed random sampling based on the prior distributions defined in Table 1 and generated MBHB signals for testing.
Table 3: In the 850 simulations, we compared the correlation coefficients between the true values and the recovered values for the four models. In addition, in the last two columns of the table, we list the network parameters and training times for the three models.
$\mathcal{M}_{c}$ $q$ $\chi_{z1}$ $\chi_{z2}$ $d_{L}$ Network parameters Training time
CNFs VPFMPE-TResMlp 0.9978 0.9790 0.9723 0.8163 0.9408 $\approx$ 189 millions $\approx$ 2.0 days
FMPE-TResMlp 0.9978 0.9787 0.9719 0.8139 0.9400 $\approx$ 189 millions $\approx$ 2.0 days
FMPE-DeResNet 0.9970 0.9762 0.9563 0.7174 0.9397 $\approx$ 211 millions $\approx$ 2.3 days
DNFs NPE-DeResNet 0.9261 0.8647 0.6620 0.1681 0.8634 $\approx$ 384 millions $\approx$ 6.5 days
We injected 850 waveforms drawn from the prior shown in Table 1, along with residual instrumental and confusion noise. We will take the median of the recovered values as the point estimate and calculate the correlation coefficient with the true values of the parameters. The correlation coefficient reflects the degree of linear correlation between the true value distribution and the estimated values. A coefficient value of 1 indicates perfect agreement between the estimated and true values. Note that the reference times of these testing signals are randomly distributed throughout a year, rather than being fixed to the 1st day.
Our results are presented in Table 3. In the final column of the Table 3, we have included the approximate training times for the four models. For $t_{c}$, $\varphi_{c}$, $\iota$, $\beta$, $\lambda$, and $\psi$, due to the presence of multimodality (i.e. there is more than one peak in the posterior distributions), using the median as a point estimate is not applicable. Our results indicate that FM with CNFs reduced training time by a factor of three and achieved complete unbiased estimation for the injected signals. In contrast, NPE with DNF not only had a longer training time but also produced significant distortions in $\chi_{z1}$ and $\chi_{z2}$, potentially leading to erroneous scientific analysis. Among the methods we tested CNF for MBHBs, the VPFMPE-TResMlp model achieved the best results and exhibited the fastest training time.
@@ 4 Experiments
Based on the analysis of P-P plot (Figure 5), we confirmed that the model can achieve unbiased estimates when considering only the instrumental noises. However, the confusion noise, which is composed of the GW signals from tens of millions of GBs, is also one of the main noise sources in the mHz frequency band. Currently, an accurate GB population model is still absent, and the characterization of confusion noise is typically conducted after the identification of bright signals such as MBHBs in the whole global fitting pipeline. Considering these practical factors, we have not included confusion noise in the training set. Instead, in order to assess the model’s applicability under real detection scenarios, a simulated confusion noise is incorporated in the test set.
To date, existing studies have not been successful in using likelihood-free inference to provide unbiased estimates of the MBHB across the entire parameter space. For example, Ref. [37] only focused on four parameters, and Ref. [32] have limited to positive spins. Also, we use FM with CNFs for complete parameter estimation of MBHBs and achieve comparable results to nested sampling. Therefore, in subsequent comparisons, we will only compare the results of VPFMPE-TResMlp to those of the MCMC method.
@@
The aim of using CNFs for MBHBs is to reproduce the comparable posterior distribution as the MCMC method. Thus, by comparing the results of Bayesian inference on the same event, the performance of the model can be better assessed. In this case, we use Nested Sampling [74, 75, 76] as a Monte Carlo technique for comparison experiments, as shown in Figure 6.
An example using the same prior probabilities outlined in Table 1 is provided in Figure 6, where a test injection is subjected to confusion noise on the 30th reference day. The injected GW signal originates from an MBHB with a total SNR of 1543, and its parameters are detailed in Table 4. Our model exhibited results similar to the traditional nested sampling, demonstrating consistency in the sense of the degeneracy among extrinsic parameters. This consistency aligns with the crucial physical characteristics of MBHBs. Note that compared to nested sampling, our machine learning process always produces similar posterior distributions on $t_{c}$, $d_{L}$, $\iota$, $\lambda$, and $\psi$, as shown in Figure 6. We have achieved accurate estimates of the sky position parameters. This is most advantageous since accurate and reliable estimates of the sky position parameters (longitude, latitude) are useful in guiding the search for electromagnetic observatories.
@@
Specifically, under the influence of confusion noise, we have collated the true values of the injections alongside the results generated by both the model and the nested sampling process in Table  4. The test results show that, except for the parameters $\beta$, $\chi_{z1}$ and $\chi_{z2}$, our machine learning model is consistent with the posterior distributions of nested sampling in terms of order of magnitude. However, we also observed a decrease in the model’s performance in estimating $\beta$, $\chi_{z1}$ and $\chi_{z2}$ under the influence of confusion noise, but they are generally still within the 2 $\sigma$ or 3 $\sigma$ range. Such accuracy of estimation is basically sufficient for a rapid preliminary parameter estimation, and manifests the robustness of our model when applied to data outside the training set. Addressing the issue of confusion noise will be regarded as a key focus of our future work, in order to further enhance the robustness and accuracy of the model.
In the final part of the results, we explore the computational efficiency of the algorithm. Table 5 shows the speed of generating posterior samples using nested sampling compared to the VPFMPE method. It is important to note that for nested sampling the search is performed near the near the true intrinsic parameter values of the injected MBHB. In practice, the time consumed by nested sampling is significantly longer than the times listed in the table. Even under these conditions, our VPFMPE method remains two orders of magnitude faster than the nested sampling method. Additionally, the VPFMPE method employed an NVIDIA A800 GPU for sampling, while nested sampling utilized an Intel(R) Xeon(R) Platinum 8268 CPU.
@@ 5 Conclusions
In this study, we have demonstrated the inaugural application of using CNFs for MBHBs in the field of space-based GW detection, taking the parameter estimation of merging MBHBs as a representative example. We recommend the FM with CNFs, particularly the VPFMPE-TResMlp method, as a promising approach for the inference of merging MBHBs.
It is shown that our method have achieved comprehensive statistical inference for the 11-dimensional parameters of MBHBs. This approach, as opposed to previous AI-based inference methods, is featured by faster training and more accurate estimation.
Besides, we propose the integration of model generation strategy based on the symmetry of the system with CNFs. This enables us to simplify the training task, while still obtaining a model that is applicable to the entire mission period. Note that the coordinate transformation and CNFs are theoretically applicable to other space-based missions such as LISA.
It is also noteworthy that our method exhibits robustness against challenges such as confusion noise and the complexities of time-varying response function. The significant speed of our method, surpassing traditional techniques by orders of magnitude, positions it as a potential tool for the preprocessing of global fitting. Moreover, the precise and reliable estimation of sky location parameters (longitude, latitude) might be beneficial in guiding the search for electromagnetic observatories.
Overall, the acceleration in data processing and analysis not only streamlines workflows but also broadens the horizons for gravitational wave data exploration. These advancements herald a transformative era in space-based GW detection, fortifying the capabilities of missions like Taiji and bringing us closer to unraveling the mysteries of the universe.
Furthermore, the application of CNFs to other signals remains to be explored. In this paper, CNFs are successfully applied to short-duration MBHB merger signals. However, the following two factors may pose challenges for the further applications of CNFs on other signals. Firstly, EMRIs and GBs typically require months of observation time to achieve considerable SNRs, hence significantly increasing the dimensionality of data. Secondly, EMRIs are known for their complicated waveforms and multiple local maxima on the likelihood surface [77].
6 Data access statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.